# Beyond Exascale: Dataflow Domain Translation on a Cerebras Cluster

### v1

@@ 3. Implementation on Wafer Scale Engine
This section describes the domain translation software framework that we developed and the hardware used in our experiments.
@@ 3.1. Stencil Implementation
We implemented a generic stencil-code framework in the Tungsten dataflow language (Santos et al., 2024). The framework allows declaring state variables, defining time-stepping operators, and invoking translation primitives to moves data by a $(p,p)$-step in the grid space with a combination of memory copy and network transmission.
@@
3xp param n,p; // Grid size n x n, and p = stencil width = translation distance
4xp const w = 2*p; // Width of communication layer
5sp x[n+w][n+w]; // Solution
6sp y[n+w][n+w]; // Temporary to hold next timestep
7sp avec[5]; // Stencil weights: a(-1,0), a(0,-1), a(1,0), a(0,1), a(0,0)
8 // a(*,*) set to reflect a 5-point Laplacian in space with
9 // forward Euler intgeration in time.
10
11function sendrecv() {
@@
[…]
21 }
22}
23
24function compute() {
25 let i $\in$ [0,n);
26 let j $\in$ [0,n);
27 sp a;
28 a = avec[0];
[…]
30 a = avec[1];
[…]
32 a = avec[2];
[…]
34 a = avec[3];
[…]
36 a = avec[4];
[…]
38
@@
Each core holds $n\times n$ grid points. At compile time, we specify $n$; $(w_{\mathrm{w}},w_{\mathrm{h}})$, the number of cores to use on each node; and $q$, the amount of node-to-node interconnection bandwidth to allocate. To load the program on a cluster, we provide a graph of nodes (identified by their network addresses) for establishing software-defined network links that connect adjacent subdomains. The graph has annotations with initial condition data for each subdomain.
The framework initiates program execution by setting all nodes to run. This is an inherently an asynchronous process. For consistent performance measurements, we synchronize the wafers and cores before beginning the time stepping loop.
@@ 3.3. The Wafer-Scale Engine Cluster
A wafer-scale engine (WSE) (Rocki et al., 2020) is a system built with the largest processor chip ever produced, and is an example of a spatial architecture. A WSE comprises a 2D grid of tiles, each composed of a processing element (PE), local memory, and a router. Each router has five ports, connecting one each to its immediate cardinal neighbors and to the local PE. Routers forward 32-bit messages called wavelets with very low, order nanosecond latency. Each router has 24 virtual channels. A virtual channel multicasts traffic arriving from any one of the five ports to any subset of the five ports. Routers can be dynamically reconfigured in response to specially formatted wavelets.
@@
Figure 4. Switch Topology for an arbitrary number of WSEs. The same port index for all wafers in a cluster live on the same network switch. :
@@
The node-to-node network links including the order 2,000 tile network hops and ethernet switch to bridge vertically adjacent subdomains has a routing latency of about $10\mathrm{\mu{s}}$. The average link latency for using the sequential node performance model (Section 4.1) is about $1000^{-1}(10\mathrm{\mu{s}}+1000\mathrm{ns})=1.1\mathrm{ns}$. The domain translation algorithm allows us to benefit from the low average link latency whereas ghost methods are limited by the worst-case $10\mathrm{\mu{s}}$ latency.
@@
Figure 5. An example cluster topology showing the original spatial stencil and its Y-axis mirror which alternate between every wafer in the cluster. Mirroring enables data to enter and exit through the same switch, minimizing communication latency.:
Figure 6. On-wafer routing. Horizontal flow (purple) enters the node’s left edge and creates a daisy chain between every pair of horizontally adjacent cores. It egresses the system’s right edge. Vertical flow (dark green) enters from the right edge and makes a turn toward the bottom edge. The bottom row of cores receive this message and retransmit it upwards in a daisy chain (light green) among every pair of vertically adjacent cores. The top row of cores transmit the message back down (blue) with a turn to egress the node at its left edge. This scheme uses 8 virtual channels. Both daisy chains use two virtual channels alternating channels on an even/odd scheme. Both turns use two virtual channels with the change occurring at the bend.:
@@ 4. Performance Characterization
This section describes the specific experiments we ran for the close match between the overfed and predicted performance for the implementation.
@@ 4.1. Methodology
To assess scaling and fit a performance model, we ran a sweep of performance tests varying weak scale, strong scale, and node performance (Table 5). At regular timestep intervals, all cores record their 48-bit hardware clock-cycle counter to local memory. Instrumented runs produce an array with dimensions $(\mathrm{cluster_{w}}\times\mathrm{cluster_{h}}\times\mathrm{ncore_{w}}\times\mathrm{ncore_{h}}\times\mathrm{samples})$. For large runs (50M cores, 1000 samples) this produces $300~\mathrm{GB}$ of telemetry. Therefore we only retain data from a subset of cores for analysis.
@@
We fit models to three different codes: the heat equation with 5- and 9-point stencils, and the Shallow Water equations. The compute cost of the performance models are shown in Table 6. In the limit of large $n$, all model become compute bound, and the asymptotic utilization is shown in the table.
Code Performance model Flops/step Asymptotic utilization
@@
SWE $f(n)=1026+183.2n+137.6n^{2}$ $155n^{2}$ 56%
Table 6. Performance models for different PDE simulations. The models were fit to single node experiments not using external I/O. $f(n)$ gives the estimated cost of one timestep in clock cycles. $n$ indicate the problem size, where the number of grid points per core is $n\times n$. The utilization is measured in fraction of peak flops/s (2/cycle)
We characterize a node’s compute-bound performance in an IO-unconstrained environment. Here, on-wafer routing directly bridges opposite edges of the core array. We regress iteration time’s dependence on grid points per core as $t=\sum_{i=0}^{d}a_{i}n^{i}$.
IO-bound performance is assessed by measuring link payload per iteration. We independently regress horizontal and vertical payloads as $(b_{0},b_{1},b_{2})\cdot(nw_{\mathrm{w}},w_{\mathrm{w}},1)$ and $(d_{0},d_{1},d_{2})\cdot(nw_{\mathrm{h}},w_{\mathrm{h}},1)$. The current implementation has two IO limitations:
@@ 4.2. Measured results
Using the heat equation as our baseline experiment, we conducted a range of experiments on a cluster of wafer-scale CS-3 systems (Wang, 2024). We varied the number of nodes and the number of grid points per core (Table 5). We used the resulting measurements to assess the weak scaling of the performance measured in time steps per second. While varying the number of CS-3 nodes from 4 to 64, we observed near-perfect weak scaling with a remarkably consistent performance for each setting of grid point number per core (Figure 7).
@@
Figure 7. Weak scaling of the heat equation for runs on 4 to 60 CS-3 node counts. The observed performance was remarkably consistent for each points/core setting. :
@@
Figure 8. Comparison of model prediction with measured results. The model correctly shows the cross-over between IO-bound and compute-bound behavior.:
Having established a strong alignment of model predictions with measured data, we can plot strong scaling. The left panel of Figure 9 shows our performance model, including I/O, for scaling problems of various sizes across 4 to 128 CS-3 nodes, as well as data points measured on cluster simulations. We see close agreement between our model and real measured performance. Our systems allow boosting the clock frequency from the standard 750MHz up to 1.2GHz. At this operating point, the system is power constrained and can throttle significantly. We created a power optimzed version of the 9-point heat equation that takes advantage of a small cache of power efficient memory in each core. The right panel of Figure 9 weak scaling performance for this code operating at 1.2GHz with 3 billion grid points per node. The weak scaling behavior is nearly perfect and we achieved 84.7 PFLOPS using 64 nodes. We plan to extend our work to larger clusters and expect to observe increased performance consistent with model predictions.
The power specification per CS-3 node is 23kW. Thus, the power-efficient performance of the 64-node CS-3 cluster measures at 57 GFLOPs per Watt. In comparison, the current leader of Green500 (JEDI) achieves 72.7 GFLOPs per Watt on a dense linear algebra workload (Top500, 2024). To the best of our knowledge, there is no sparse computation on any cluster that is near the efficiency we observe with the CS-3 cluster.
Figure 9. (left) Strong scaling for the 5-point heat equation code on different number of WSE nodes, configured at the standard 750 MHz clock frequency. It shows perfect scaling up to 64 nodes for the largest simulations, and very close agreement with the performance model. (right) Weak scaling for 9-point heat equation on clock-boosted system. The code was specially optimized to reduce power throttling at the higher clock frequency. On 64 nodes and 192 billion grid points, we achieved 84.7 PFlops.:
We tested the domain translation method for shallow water equations using the GEBCO_2024 Grid (Group, 2024). The data set offers a global continuous terrain model for ocean and land with a spatial resolution of 15 arc seconds or 462 meters.
@@
As with the heat equation, we observed near-perfect weak scaling across multiple CS-3 nodes. We also measured more than one sustained flop per cycle in the compute-bound regime.
Figure 10. Planetary-scale tsunami wave propagation 14 hours after asteroid impact with energy of 2.4M tons of TNT equivalent.:
Figure 11. Wave impact at San Francisco Bay:
@@ 5. Conclusions
For these reasons, we were careful to demonstrate a meaningful result for science at planetary scale. The shallow water equations solved here represent the core of several important geophysical applications. In addition to the tsunami application described above, shallow water solvers are critical components of both atmospheric and oceanic modeling. In these disciplines, one first constructs a shallow water model to test numerics and computational implementation, and then extends these models to three-dimensional atmosphere and ocean models used for numerical weather prediction and earth system modeling. For example, the state-of-the-art CESM, E3SM, FV3 and MPAS atmosphere and ocean models ((Skamarock et al., 2012; Ringler et al., 2013; Small et al., 2014; Golaz et al., 2022; Dahm et al., 2023; Donahue et al., 2024) are derived from the shallow water models given in (Putman and Lin, 2007; Thuburn et al., 2009; Taylor and Fournier, 2010)).
@@
Our results running shallow water equations on the Cerebrus system demonstrate the promise of this architecture for atmosphere and ocean models critical for weather prediction and earth systems assessments. These models traditionally run at less than 5% of peak (Carman et al., 2017). Extending our shallow water approach to a full model running at 60% of peak would thus represent a breakthrough in performance per Watt and also in terms of total achievable throughput.

### v2

@@ 4. Implementation on Wafer Scale Engine
This section describes the domain translation software framework and the hardware used in our experiments.
@@ 4.1. Stencil Implementation
We implemented a generic stencil-code framework in the Tungsten dataflow language (santos2024breaking). The framework allows declaring state variables, defining time-stepping operators, and using translation primitives to move data by a $(p,p)$-step in the grid space with a combination of memory copy and network transmission.
@@
The code is parametrized by the number grid points ($n\times n$), the compute grid dimensions $(w_{\mathrm{w}},w_{\mathrm{h}})$, and the amount of node-to-node interconnection bandwidth ($q$) to allocate. We specify these parameters at compile time. To load the program on a cluster, we provide a graph of nodes (identified by their network addresses) for establishing software-defined network links that connect adjacent subdomains. The graph has annotations with initial condition data for each subdomain.
The framework initiates program execution by setting all nodes to run. This is an inherently asynchronous process. For consistent performance measurements, we synchronize the wafers and cores before beginning the time stepping loop.
@@
3// Grid size n x n
4// p = stencil width = translation distance
5xp param n,p;
6xp const w = 2*p; // Width of communication layer
7sp x[n+w][n+w]; // Solution
8sp y[n+w][n+w]; // Temporary to hold next timestep
9
10// Stencil weights: 5-point Laplacian in space,
11// forward Euler integration in time
12sp avec[5]; // a(-1,0), a(0,-1), a(1,0), a(0,1), a(0,0)
13
14function sendrecv() {
@@
16 // Send right and up
[…]
19 // Receive from left and from below
[…]
22 }
23 parallel {
24 // Send corner data received from left to above
[…]
26 // Receive corner data from below
[…]
28 }
29}
30
31function compute() {
32 let i $\in$ [0,n);
33 let j $\in$ [0,n);
@@ 4.3. The Wafer-Scale Engine Cluster
A wafer-scale engine (WSE) (WSE_HPC_2) is a system built with the largest processor chip ever produced, and is an example of a spatial architecture. A WSE comprises a 2D grid of tiles, each composed of a processing element (PE), local memory, and a router. Each router has five ports, connecting one each to its immediate cardinal neighbors with less than 2 ns latency, and one to the local PE. Routers forward 32-bit messages called wavelets on 24 virtual channels. A virtual channel multicasts traffic arriving from any one of the five ports to any subset of the five ports. Routers can be dynamically reconfigured in response to specially formatted control wavelets.
@@
Figure 4. Switch Topology for an arbitrary number of WSEs. The same port index for all wafers in a cluster live on the same network switch.
High-level View of WSE
@@
At the largest compute grid size $w_{h}\sim w_{w}\sim 1,000$, the node-to-node transmission involves up to 1,000 fabric hops on the wafer to get to the edge, an ethernet link through a switch to the next wafer, and up to 1,000 fabric hops to get to the receiving tile. The ethernet routing latency is close to $10\mathrm{\mu{s}}$ and the average link latency for using the sequential node performance model (Section 5.1) is about $1000^{-1}(10\mathrm{\mu{s}}+2000\mathrm{ns})=12\mathrm{ns}$. The domain translation algorithm allows us to benefit from the low average link latency whereas ghost methods are limited by the worst-case $10\mathrm{\mu{s}}$ latency.
@@
Figure 5. An example cluster topology showing the original spatial stencil and its Y-axis mirror which alternate between every wafer in the cluster. Mirroring enables data to enter and exit through the same switch, minimizing communication latency.
Cluster connectivity.
Figure 6. On-wafer routing. Horizontal flow (purple) enters the node’s left edge and creates a daisy chain between every pair of horizontally adjacent cores. It egresses the system’s right edge. Vertical flow (dark green) enters from the right edge and makes a turn toward the bottom edge. The bottom row of cores receive this message and retransmit it upwards in a daisy chain (light green) among every pair of vertically adjacent cores. The top row of cores transmit the message back down (blue) with a turn to egress the node at its left edge. This scheme uses 8 virtual channels. Both daisy chains use two virtual channels alternating channels on an even/odd scheme. Both turns use two virtual channels with the change occurring at the bend.
On-wafer routing.
@@ 5. Performance Characterization
This section describes our performance modeling and the specific experiments we ran. As we will show, the models closely match the experiments at all scales.
@@ 5.1. Methodology
To assess scaling and fit a performance model, we ran a sweep of performance tests varying weak scale, strong scale, and node performance (Table 5). At regular timestep intervals, all cores record their 48-bit hardware clock-cycle counter to local memory. Instrumented runs produce an array with dimensions $(\mathrm{cluster_{w}}\times\mathrm{cluster_{h}}\times\mathrm{w_{w}}\times\mathrm{w_{h}}\times\mathrm{samples})$. For large runs (50M cores, 1000 samples) this produces $300~\mathrm{GB}$ of telemetry. Therefore we only retain data from a subset of cores for analysis.
@@
We fit models to three different codes: the heat equation with 5- and 9-point stencils, and the Shallow Water equations. The compute cost of the performance models are shown in Table 6. In the limit of large $n$, all models become compute bound, and the asymptotic utilization is shown in the table.
Code Performance model Flops/step Asymptotic utilization ($n\rightarrow\infty$) Realized utlization I/O (words/step)
@@
SWE $f(n)=1026+183.2n+137.6n^{2}$ $155n^{2}$ 56% 53% $n=24$ $64n+80$
Table 6. Performance models for different PDE simulations. The models were fit to single node experiments not using external I/O. $f(n)$ gives the estimated cost of one timestep in clock cycles. $n$ indicate the problem size, where the number of grid points per core is $n\times n$. The model error is $5\%$ for $n<=5$, and $0.3\%$ for $n>5$. The utilization is measured in fraction of peak flops/s (2/cycle). The I/O columns lists how many floating points words are received each timestep.
We characterize a node’s compute-bound performance in an IO-unconstrained environment. Here, on-wafer routing directly bridges opposite edges of the core array. We regress iteration time’s dependence on grid points per core as $t=\sum_{i=0}^{d}a_{i}n^{i}$. The resulting models are presented in Table 6 and are accurate to a fraction of a percent for $n>5$.
IO-bound performance is assessed by measuring link payload per iteration. We independently regress horizontal and vertical payloads as $(b_{0},b_{1},b_{2})\cdot(nw_{\mathrm{w}},w_{\mathrm{w}},1)$ and $(d_{0},d_{1},d_{2})\cdot(nw_{\mathrm{h}},w_{\mathrm{h}},1)$. The I/O model assumes that the left and right side of the wafer each has 300 Gbit/s network bandwidth in each direction. Table 6 lists the amount of I/O between a core and its neighbors for each code.
The current implementation has two IO limitations:
@@
In the strong scaling limit (small $n$) these limitations can have significant impact, and resolving both of them could improve timstepping rate by up to 50% for $n=2$.
@@ 5.2. Measured results
5.2.1. Characterization of 5-point heat equation
The 5 point heat equation is the most data-intensive. For that reason, we choose it for our main measurement and characterization. We conducted a range of experiments on a cluster of wafer-scale CS-3 systems (CS3) running at a 750MHz clock frequency. We varied the number of nodes and the number of grid points per core (Table 5). We used the resulting measurements to assess the weak scaling of the performance measured in time steps per second. While varying the number of CS-3 nodes from 4 to 64, we observed near-perfect weak scaling with a remarkably consistent performance for each setting of grid point number per core (Figure 7).
@@ 5.2.1. Characterization of 5-point heat equation
Figure 7. Weak scaling of the heat equation for runs on 4 to 60 CS-3 node counts. The observed performance was remarkably consistent for each points/core setting.
Scaling.
@@
Figure 8. Comparison of model prediction with measured results. The model correctly shows the cross-over between IO-bound and compute-bound behavior.
Model.
Having established a strong alignment of model predictions with measured data, we can plot strong scaling. The top panel of Figure 9 shows our performance model, including I/O, for scaling problems of various sizes across 4 to 128 CS-3 nodes, as well as data points measured on cluster simulations. We see close agreement between our model and real measured performance.
5.2.2. 9-point heat equation and peak performance
Our systems allow configuring the clock frequency. We ran experiments at 750MHz and 1.2GHz. At the 1.2GHz operating point, the system is power constrained and can throttle significantly. We created a power optimized version of the 9-point heat equation that takes advantage of a small region of power efficient memory in each core. The bottom panel of Figure 9 weak scaling performance for this code operating at 1.2GHz with 3 billion grid points per node. The weak scaling behavior is nearly perfect and we achieved 84.7 PFLOPS using 64 nodes. The power specification per CS-3 node is 23kW. Thus, the power-efficient performance of the 64-node CS-3 cluster measures at 57 GFLOPs per Watt. In comparison, the current leader of Green500 (JEDI) achieves 72.7 GFLOPs per Watt on a dense linear algebra workload (Green500). To the best of our knowledge, there is no sparse computation on any cluster that is near the efficiency we observe with the CS-3 cluster.
At the 1.2GHz operating point, even the power optimized code is affected by power throttling. To further push the performance, we ran the 9-point heat equation code on a special node outfitted with an enhanced power supply. This allowed continuous unconstrained operation at 1.2GHz without power throttling. In this setting the code performed over 2.1 GFLOPS per core, and at 88% of peak performance. At 64 nodes of scale, this performance would produce 112 PFLOPS given our perfect weak scaling.
Figure 9. Top: strong scaling for the 5-point heat equation code on different number of WSE nodes, configured at a 750 MHz clock frequency. It shows perfect scaling up to 64 nodes for the largest simulations, and very close agreement with the performance model. Bottom: weak scaling for 9-point heat equation on system at 1.2GHz. Blue line: code optimized to reduce power throttling at the higher clock frequency. On 64 nodes and 192 billion grid points, we achieved 84.7 PFlops. Yellow line: system with enhanced power supply.
Strong scaling
5.2.3. The Shallow Water Equations
The heat equation is a simplistic model. To explore the efficacy of the Domain Translation method on more complex and realistic equations, we tested the domain translation method for the shallow water equations using the GEBCO_2024 Grid (gebco). The data set offers a global continuous terrain model for ocean and land with a spatial resolution of 15 arc seconds or 462 meters.
@@ 5.2.3. The Shallow Water Equations
The performance model for the SWE is shown in Table 6. As with the heat equation, we observed near-perfect weak scaling across multiple CS-3 nodes. In the compute bound regime we achieve 53% of peak performance. We plan to extend our work to larger clusters and future generations of our hardware, and expect to observe increased performance consistent with model predictions.
Figure 10. Planetary-scale tsunami wave propagation 14 hours after asteroid impact with energy of 2.4M tons of TNT equivalent.
Large-scale flow
Figure 11. Wave impact at San Francisco Bay
Large-scale flow
@@ 6. Conclusions
The Domain Translation method can hide latencies that are quadratic in the amount of memory (given by Eq. 1). In this context, an entire cluster can be considered a unit, and its aggregate memory used to cover latency. Besides the strong scaling benefits mentioned above, a further implication of our work is that clusters in different cities could be fruitfully networked and use Domain Translation to overcome the millisecond or longer network latencies, opening the door for running parallel applications across multiple exa-scale machines.
For these reasons, we were careful to demonstrate a meaningful result for science at planetary scale. The shallow water equations solved here represent the core of several important geophysical applications. In addition to the tsunami application described above, shallow water solvers are critical components of both atmospheric and oceanic modeling. In these disciplines, one first constructs a shallow water model to test numerics and computational implementation, and then extends these models to three-dimensional atmosphere and ocean models used for numerical weather prediction and Earth system modeling. For example, the state-of-the-art CESM, E3SM, FV3 and MPAS atmosphere and ocean models ((Skamarock2012; Ringler2013; Small2014; Golaz_et_al:2022; Dahm2023; Donahue_et_al:2024) are derived from the shallow water models given in (Putman2007; thuburn2009numerical; TayFour10)).
@@
Our results running shallow water equations on the Cerebras system demonstrate the promise of this architecture for atmosphere and ocean models critical for weather prediction and Earth systems assessments. This sets the stage for greatly enhanced Earth system modeling with an order of magnitude increased throughput and 1.5 order of magnitude improved power efficiency.
Acknowledgements.
The authors would like to thank Amirali Sharifian and Milad Hakimi for their work on improving the compiler to support this project. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia LLC, a wholly owned subsidiary of Honeywell International Inc. for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research’s Computer Science Portfolios program. This work was also funded by the United States Department of Energy, National Energy Technology Laboratory. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.